The mixing department has​ 23,000 units and​ $50,000 in costs for which to account. Of the​ 23,000 units,​ 10,000 were completed and transferred to the next department. The​ 13,000 remaining were​ 25% complete for conversion costs. Direct materials are added at the beginning of the​ process, and the conversion costs are added evenly throughout the process. The cost per equivalent unit is​ $3.25 for direct materials and​ $0.80 for conversion costs. The total costs to account for are?
Answer #1

Cost of goods completed And transferred out = (3.25+.80)*10000 = 40500

Cost of ending work in process = (13000*3.25+13000*25%*.80) = 44850

Total cost to account for = 40500+44850 = 85350